### v1

# The cell as a token: high-dimensional geometry in language models and cell embeddings

Abstract
Single-cell sequencing technology maps cells to a high-dimensional space encoding their internal activity. This process mirrors parallel developments in machine learning, where large language models ingest unstructured text by converting words into discrete tokens embedded within a high-dimensional vector space. This perspective explores how advances in understanding the structure of language embeddings can inform ongoing efforts to analyze and visualize single cell datasets. We discuss how the context of tokens influences the geometry of embedding space, and the role of low-dimensional manifolds in shaping this space’s robustness and interpretability. We highlight new developments in language modeling, such as interpretability probes and in-context reasoning, that can inform future efforts to construct and consolidate cell atlases.
The implicit goal of modern single-cell technologies is to decompile the cell—to abstract it away from its squishy context, and to render it as a single point in a high-dimensional vector space. But how do we know if this space is meaningful? If a single-cell embedding exactly matches known regulatory and developmental mechanisms, an embedding space may be accurate yet uninformative for making new discoveries. Conversely, if this space fails to recapitulate known relationships, it is difficult to attribute this to novelty or inaccuracy. A similar problem arises in statistical learning. Large language models are trained on vast, unannotated volumes of text. To represent diverse text sources consistently, these models initially split input text into discrete tokens, minimal units comprising words or word fragments. They then convert these tokens into vectors in a high-dimensional space, a representation that enables further processing by modern, continuously-valued learning models.
The success of language models stems, in part, from the ability of language embeddings to accurately encode syntactic and semantic structure in high-dimensional spaces. The unique properties of high-dimensional geometry allow embeddings to effectively encode the semantic structure of language along low-dimensional manifolds, mirroring findings from single-cell biology in identifying developmental pathways and rare cell types. What insights can single-cell embeddings gain from the statistical learning community? This perspective seeks to identify commonalities between these fields, highlighting general principles of language embeddings that may inform ongoing work in single-cell genomics.
.1 Context shapes the geometry of embeddings
Figure 1: Low-rank structure in high-dimensional embeddings. (A) An embedding of the full text of the novel Blood Meridian using a word2vec model originally trained on a dataset of $10^{11}$ words drawn from Google News articles [1]. Vectors are clustered using K-means partitioning, and then summarized into metagroups with a topic embedding model (colors and annotations). (B) An embedding of $6\times 10^{4}$ human peripheral blood mononuclear cells based on single-cell RNA sequencing of $1.6\times 10^{4}$ genes. Colors correspond to immune cell subtypes, as determined by marker genes for characteristic cell surface proteins like CD4, CD8, etc.
Word embeddings are often trained using variants of context masking. Given a sequence of words, “The parliamentarian led the assembly.”, a single word is masked and a supervised model is trained to fill in the blank: “The [TOK] lead the assembly.” When trained at scale, models learn to group certain words (e.g. “leader,” “president,” “speaker”) that appear in similar contexts. Theoretical motivation for this approach comes from the distributional hypothesis, which equates distances between vectors in embedding space with distances between distributions of co-occuring tokens within the training corpus [2, 3]. An early context-based embedding, word2vec, was released in 2012 and maps each of $10^{7}$ distinct words to a point in a $300$-dimensional continuous vector space [4]. This approach directly optimizes an objective function motivated by the distributional hypothesis, by training an embedding that maximizes the posterior probability of word-context pairs seen in the corpus, while minimizing the probability of randomly-generated pairs. Because this posterior objective is maximized only during training on a large text corpus, word2vec represents a static embedding: when the trained model is subsequently applied to new text, any appearance of a given token always maps to the same representation.
Current single cell embeddings are both static and context-independent. After isolation, sequencing, and demultiplexing, a cell becomes a collection of RNA transcripts, which are used to create a vector approximating the transcript counts per gene per cell [5]. At this stage, a given cell’s embedding is uniquely determined by the expression levels of each gene, thus decoupling the representation of a given cell from others. Thus, in principle, single-cell embeddings do not invoke the distributional hypothesis: a cell’s embedding is an innate property, rather than a property relative to a corpus of cells. In fact, many preprocessing schemes applied to count matrices—such as batch or cell cycle correction—serve to enforce static, context-free structure on embedding space [6]. However, data reduction methods like principal components analysis for visualization, or unsupervised clustering for cell type identification, produce context-dependent representations that depend on relative differences among cells. Context-dependence also arises when multiple datasets are merged, or when end-to-end embedding models are trained across many datasets. In an analogy to language embeddings, individual genes are letters, cells are words, and tissues are chapters (Fig. 1).
Theoretical analysis of word2vec and its variants show that these methods, in practice, factorize a rectangular matrix comprising the mutual information between each token and its context [7]. Linguistic structures, such as synonym clusters, lead to low-rank structure in this matrix [8, 9], analogous to low-rank structure in single-cell count matrices associated with cell types. Low-dimensional manifolds in single-cell embeddings typically arise from highly-coordinated biological processes, such as differentiation, which exhibit predominantly deterministic dynamics. In language embeddings, low-rank structure arises due to overparametrization—the highest-rank word-context matrix would simply represent an isotropic Gaussian distribution. Anisotropy in high-dimensional embeddings thus indicates structure in the underlying generative process, whether linguistic or biological.
A key limitation of static language embeddings stems from polysemy, in which the same token has multiple meanings [10, 11]. For example, “bank” may refer to the side of a river bed, or to a financial institution. Static word embeddings place polysemous tokens at intermediate positions in embedding space, between positions associated with their divergent meanings. Such compromises distort and curl embedding space, reducing the space’s ability to represent large-scale structure by making distances between vectors less meaningful [12, 13, 14]. As a result, static embedding models tend to underestimate differences among strongly-distinct concepts, limiting their ability to recognize hierarchies among words [15].
In gene expression, curvature due to polysemy may be biological, rather than artifactual. Cellular differentiation datasets exhibit extended low curvature regions associated with stereotyped cell states, punctuated by high-curvature regions associated with transitions [16]. These transition states, such as differentiating stem cells, necessarily occupy intermediate locations in embedding space. In large-scale cell atlases, biological polysemy may arise due to pleiotropy, in which a single gene is involved in multiple biological processes [17]. However, spurious polysemy can arise due to technical errors, leading to unresolved cell subtypes or tissue groups. This effect becomes more pronounced at low read depth, resulting in missing genes or greater sampling error in counts. For example, blood vascular endothelial cells share relatively consistent transcriptional profiles, due to their similar structural roles across different tissues [18]. Endothelial cells from different tissues thus tend to map to the same area in embedding space, despite their anatomical separation. Resolving this ambiguity either requires additional marker genes and greater sequencing resolution, or assays that barcode transcripts with additional information. For example, cell painting produces a high-dimensional vector of morphological features extracted from fluorescent microscopy [19], while CITE-Seq augments each transcript with information about cell surface proteins—thus avoiding cellular polysemy [20].
Recent language models introduce dynamic embedding, in which a given token’s embeddings vary after training based on the context in which the token appears [21, 11, 22]. During inference, a token’s representation is combined with a vector derived from many neighboring tokens, resulting in a joint representation. Even after training is complete, a given token’s embedding will vary across different tasks. Thus, while static embeddings map each token to a single point in embedding space, dynamic embeddings map each token to a continuous manifold capturing the diverse contexts in which it appears. The distance between the same token in different contexts is smaller than the distance between tokens, consistent with low-dimensional, anisotropic structure [23].
Dynamic embedding can improve the global structure of embeddings, particularly in organism-scale cell atlases. Spatial transcriptomics pairs each transcript with information about the cell’s absolute spatial position, or relative position among neighboring cells. As a result, embeddings learned by these methods encode an underlying metric, and so both local and global distances are meaningful [24, 25]. More abstract context information, such as organ group or tissue annotations, improve embeddings by disambiguating similar transcriptional profiles arising in distinct contexts [26]. Conceptually, these approaches resemble language models that combine tokenization with queries to an external database, which provides richer context [27, 28]. This can include structured sources of information about tokens, like encyclopedias or human-curated concept maps [29, 30]. The resulting models capture global relationships among concepts, without requiring substantial additional training. For single-cell data, similar approaches enrich transcript information with tissue or preexisting cell type annotations [31, 32, 33], gold-standard experimental associations or transcription factors [34], gene ontologies [35], or even topic information from scientific literature databases [36, 37]. Other efforts pair each cell with gene-level context (such as sequence position or chromatin accessibility) to highlight mechanistic relationships [38, 39].
Advances in language and cell embeddings converge in recent single-cell foundation models, which combine massive gene expression datasets with auxiliary information to develop robust, multimodal representations [40, 41, 39, 36, 37]. Newer embedding schemes, such as contrastive learning with negative sampling, directly train models to distinguish cells appearing in similar versus different contexts [42]. This approach endows representations with an underlying similarity metric, ensuring distances (and not just neighbors) are meaningful in embedding space. As single-cell foundation models grow larger and more sophisticated, embeddings may be further enriched by augmented retrieval, a procedure by which language models query external databases of structured knowledge [43]. Augmented retrieval decouples training of representations from relevant information itself, and thus allows the structure of embeddings to update dynamically as new experimental data becomes available.
.2 Analogies as manifolds in embedding space
Figure 2: Analogies and low-dimensional manifolds. (A) Embeddings of particular sequences of tokens using the model of Fig. 1, with examples of escalating manifolds (red and blue lines), which overlap in regions with similar meaning (weak polysemy). A token with strong polysemy appears at an intermediate location (purple circle). An example of an analogy relationship encoded as nearly-congruent difference vectors (turquoise arrows). While nonlinear embedding methods like UMAP distort the local metric over large scales [44], the nearby position of the two analogy vectors’ heads and tails protects their congruency. (B) RNA Velocity applied to developing endocrine cells in the pancreas [45, 46]. Vectors correspond to development direction, and color corresponds to pseudotime assigned via diffusion components. Cell types along the differentiation axis are overlaid.
Language embeddings exhibit a surprising ability to encode semantic relationships as distances in embedding space [47]. For example, the difference vector from “Sacramento” to “California” in embedding space may match the vector from “Austin” to “Texas.” As a result, word2vec is capable of solving unseen analogy problems from college admissions exams, even without retraining (Fig. 2)[48, 49]. Analogies represent a standard test of reasoning ability, and embedding space unfolds computation into a high-dimension space in which reasoning coincides with vector arithmetic [50]. In this sense, word embeddings foreshadow modern works on in-context learning and zero-shot inference, in which sufficiently large models are able to perform tasks not seen during training [51].
Organismal cell atlases typically exhibit well-defined clusters associated with distinct tissue types. However, cells with different compositions but similar functions can nonetheless occupy similar relative locations in embedding space. For example, immune cells form subtypes within different organ groups, such as Kupffer cells in the liver or microglia in the brain. In whole-organism cell atlases, these cells typically appear in separate clusters associated with their primary organ groups. However, within each organ cluster, they occupy similar positions relative to other cells, underscoring their analogous roles [52, 53, 54].
Language embeddings can also capture continuous relationships among tokens. For example, escalating sequences like “good,” “better,” or “best” map to linear sequences in embedding space [47]. The distribution of vectors produced by word embeddings exhibits high anisotropy [55, 23], with embeddings spanning low-dimensional manifolds within the higher-dimensional representation space. Depending on the language, this manifold has effective dimensionality $\sim\!10^{1}$, even when the feature dimension is $\sim\!10^{2}$ [56]. These low-dimensional semantic manifolds capture phenomena such as gradations in meanings among similar words, shifts in a word’s meaning over time, singular-plural pairs, or groups of synonyms [57]. Consistent with these manifolds representing subspaces, the performance of embeddings in downstream tasks initially increases as embedding dimension increases, before plateauing at a fixed multiple of the manifold dimension [58].
Similar low-dimensional manifolds arise in cell embeddings. Across different datasets, cell replication cycles and circadian rhythms form rings [59], spatially-extended tissues form grids [60, 25], and cell differentiation hierarchies form branches [61]. In one well-known case, populations of blood cells of mixed maturity form a pitchfork structure in embedding space, illustrating a continuous progression from stem cells to red or white blood cells [61]. Gene expression manifolds have a typical intrinsic dimensionality $\sim 10^{1}$, compared to the $\sim 10^{5}$ genes measured in a typical single-cell experiment [16]. Just as word embeddings trace their properties to low-rank structure in the word-context mutual information matrix, biological processes confer low-rank structure on count matrices [62, 63].
Traversing embedding space maps the structure of low-dimensional manifolds. Theoretical models of word embeddings frame text generation as a stochastic dynamical system, with sentence formation as a random walk in token embedding space [64, 65, 66]. Under this framework, semantic manifolds act as kinetic traps for the walk, with synonymous tokens acting as basins, and connective phrases acting as bridges. In single-cell analysis, diffusion maps simulate the action of many random walks through expression space [67, 68]. These methods represent a standard approach to calculating pseudotime, which orders unsorted cell embeddings to identify temporal progressions of cells (like developmental stages) [69, 70]. In dynamic word embeddings, or in static embeddings trained on corpora from different historical periods, the relative positions of words gradually shift over time. This semantic drift may be quantified using a calculation resembling pseudotime [71].
Many features of natural language, such as parts-of-speech, intensifiers, and modifiers, recur across languages that otherwise exhibit very different syntax. For example, while English and Sanskrit have different character sets and inflections, they exhibit similar verb conjugations and noun declensions. Machine translation models must disentangle these distinctions in order to construct maps between different languages’ embedding spaces. In an encoder-decoder translation model, a sentence is first embedded using a language-specific model, then encoded into a universal representation within a latent space, and then decoded into the target language [72]. Syntax information, like word ordering or inflection, is typically distinct among languages and thus not necessarily preserved in the latent space. In contrast, semantic manifolds remain conserved, and the latent coordinates of these low-dimensional representation capture information such as token positions and parts-of-speech [73, 74]. Taking this approach even further, cross-lingual translation models construct a single shared latent space from many languages. These models typically outperform single-pair translation models, particularly for languages with less available training data, like Swahili or Urdu [75].
Could the same effect hold for rare cell types? One analogy for cross-lingual alignment is shared embeddings of cell types across distinct organisms. Statistical alignment methods may be used to combine cell type populations across species with similar tissue groups [76, 77, 78, 79, 80, 81]. Like understudied languages, rarer cell types benefit from integrated analysis; for example, in a joint embedding of human and mouse pancreatic cells, a combined embedding better resolves subpopulations associated with stress response during protein assembly [76]. Recent works extend this concept by proposing universal cell embeddings, in which a single foundation model is trained on data spanning subjects, species, and even sequencing modalities [82, 83]. The resulting embedding exhibits emergent properties, including informative zero-shot embedding of new species or tissues without retraining.
As single-cell methods become more contextual and high-dimensional, biological insights will require systematic exploration of the geometry of embedding spaces. In language modelling, tools from topological data analysis may be used to identify singularities in embedding spaces due to polysemy [12]. For larger models, techniques like linear probes or sparse autoencoders analyze a model’s internal representations of text. Linear probes are trained to predict particular linguistic features, like parts-of-speech or subject-verb agreement, from hidden states, thus quantifying how explicitly different features are represented [84, 85, 86]. Sparse autoencoders identify minimal subsets of features responsible for encoding linguistic features, thus highlighting compositionality in embedding space [87]. In gene expression datasets, these approaches may reveal implicit biological “grammars,” such as cell-cell interactions or indirect gene regulatory pathways.
.3 Task-independence and amortization of reasoning
Word embeddings derive their utility from their independence from downstream tasks. Training new language embeddings typically requires access to large amounts of computing resources, with contemporary models like DeepSeek-R1 optimizing as many as $10^{11}$ parameters over $10^{13}$ language tokens [88]. However, once trained, these models may be used as a preprocessing stage for downstream tasks, like sentiment classification or fake news detection. Single-cell technologies share this goal: if sequencing depth and quality are sufficient to capture all relevant information about a cell, then an experimenter is free to first measure all relevant variables, and later apply post hoc analysis to identify genes associated with pathologies or treatment responses.
Task-independent embeddings thus serve to amortize computation. Suppose an embedding space has linear structure, and thus the ability to render reasoning tasks like word analogies as vector operations. In this case, reasoning already occurred in the process of constructing the embedding space in the first place—downstream analysis proceeds through linear operations like vector arithmetic only because elaborate nonlinear transformations learned to disentangle intertwined concepts. Embeddings thus represent one motivation for the emerging foundation model paradigm in both language modelling and single-cell analysis, which argues that large-scale pretraining on diverse datasets can lead to simpler starting representations for smaller-scale tasks. Conceptually, if an embedding space robustly captures standard biological processes like the cell cycle, then unusual effects of a new treatment or perturbation should be easier to identify.
An emerging area in large-scale pretrained models is inference-time computing, in which large models spontaneously solve new tasks without additional training [51]. For example, large language models can be prompted to produce poetry with meter and scansion that are unseen in their training corpus [89]. The underlying mechanism, in-context learning, exploits the emergent ability of large models not only to retrieve, but also process, information during inference. Achieving such effects within an interpretable embedding space represents a frontier for single-cell foundation models. Several recent models exhibit forms of inference-time reasoning, such as zero-shot embedding of novel cell types, prediction of protein interactions, and anticipation of responses to genetic perturbations [90, 91]. However, it is difficult to clearly assess the difficulty of these tasks, because many informative biological pathways are highly indirect and nonlinear [92, 93]. For example, the immune system implements complex logical operations through elaborate combinatorial receptor-ligand interactions, phosphorylation cascades, and feedback loops, in order to discriminate self from non-self antigens [94, 95]. Parsing these logical circuits is akin to solving a complex mathematical reasoning problem, requiring models that can effectively process symbolic information. In language models, inference-time symbolic reasoning emerges at sufficient scale, with models recently advancing from solving elementary-school word problems to undergraduate mathematics [96, 97, 98]. As single-cell embedding models grow in scale, their ability to identify increasingly intricate regulatory logic will represent a key test of their generalization.
I Acknowledgments
W.G. thanks the members of the Chan-Zuckerberg theory in biology group. This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation. W.G. is supported by NSF DMS 2436233 and NSF CMMI 2440490, as well as Medici Therapeutics.